OpenStudy (anonymous):

Find∫ 0 up to 4 of (9-3x^2)dx The only part I can figure out is a=0 and b=4. I believe I have to use sigma notation.

OpenStudy (anonymous):

\[\int\limits_{0}^{4}(9-3x^{2})dx = [9(4-0)]-3*[1/3(4^{3}-0^{3})]= 9*4 - 4^{3} = 36 -64 \] I hope its correct

OpenStudy (anonymous):

Yes it is thank you so much that thing has been so hard to figure out!

OpenStudy (anonymous):

welcome :D